(1 point) Let f(x) = (x² + 5) lue). Find f (a). f(x) = Preview My Answers Submit Answers You have attempted this problem 0
Answer #1

A y= f(x) = (22+5) y= f(x) = dy Notation da Taking un on both sides ܠܫܟ݂ܛܛܛ In y = In (c²ts) In (oc). In (x²+5) eny differe